# Pseudo-Bifurcations in Stochastic Non-Normal Systems

### v2

@@ Abstract
Many of the dynamics driving complex systems in nature and society are marked by abrupt changes that result in transitions to fundamentally different behaviors [1, 2, 3, 4]. In recent decades, the dominant framework for understanding these transitions has centered around the concept of bifurcations [5, 6, 4, 7]. A bifurcation is a point at which a small change in a system’s control parameter causes a qualitative shift in its equilibrium state, resulting in a new and distinct configuration of system behavior. There are numerous examples of systems that display dynamics resembling those of bifurcations. For instance in ecology, predator-prey dynamics can lead to the sudden extinctions of a species [8], small shifts in climate patterns may induce inter-glacial periods [9], vegetation losses above a critical threshold results in desertification [10], or nutrient overload can cause a lake to suddenly transition from a clear-water to a algae dominated state [11]. In social systems, notable examples are financial markets experiencing long periods of low volatility followed by abrupt crashes [12] and growing dissatisfaction with political conditions, which can trigger widespread unrest or revolution [13]. Health-related examples involve the sudden onset of conditions like microbiome dysregulation [14], transitions between different states representing distinct illnesses based on varying internal and external influences [15], depressive episodes [16], or epileptic seizures [17].
@@
As a system approaches a critical bifurcation threshold, it is often marked by increased variance and autocorrelation in observable metrics. These statistical signatures reveal that the system’s recovery from perturbations is slowing as it approaches bifurcation, alongside an increasing sensitivity to fluctuations in the system’s state–two complementary indicators that signal the proximity to a critical transition. It has thus been suggested that these statistical signatures serve as important early-warning signals for the onset of a looming bifurcation point [5, 6, 4, 7].
@@
Despite the insights offered by bifurcation theory, the framework itself encounters at least two major challenges. First, it relies on the existence of a well-defined control parameter and its closeness to a critical threshold. Second, it assumes the presence of a driving force that pushes the system toward or beyond this threshold. These assumptions are not only challenging to verify empirically but also conceptually restrictive and lack general applicability. It is, therefore, not surprising that the question of whether certain systems are undergoing critical bifurcations continues to be a subject of debate. For example, evidence is mixed regarding whether the onset of epilepsy is truly driven by critical transitions [18, 17]. Similarly, it remains uncertain whether components of the Earth’s system, such as the Amazon rainforest or the Atlantic Meridional Overturning Circulation (AMOC), are approaching bifurcation points [19].
@@
In this work, we present an alternative, more general mechanism that relies solely on a stable equilibrium attractor with linear non-normal stochastic dynamics around it. We demonstrate that small perturbations around the equilibrium produce transient repulsions - we dub them pseudo bifurcations in this context - with similar statistical signatures to those observed near genuine bifurcations. Much like the effective dimension reduction seen during the onset of a bifurcation, we find that non-normal transients are confined to movements within a two-dimensional subspace. We further find that these transients are accompanied by statistical markers such as critical slowing down and increased autocorrelation. We empirically illustrate our findings by analyzing EEG signals that capture brain activity during epileptic seizures, revealing clear evidence of non-normal dynamics. We further demonstrate that the approach of actual bifurcations is accompanied by non-normal transients, making it difficult to distinguish between the two phenomena and potentially creating a bias that suggests the system is much closer to a critical point than it actually is. Finally, we show that the onset of non-normal transients can be anticipated by measuring the system’s alignment along its non-normal mode, and provide a numerical algorithm for stable identification of these non-normal modes. This opens a new path towards the anticipation and control of instabilities in complex systems.
@@
The only requirement to observe such transients is that the system exhibits sufficient non-normality - meaning its components interact within asymmetric hierarchical structures, a feature common in many social and natural systems [20, 21]. This characteristic is also found in non-network systems, such as turbulent flows, where non-normal dynamics arise from the interplay of viscous and advective terms, which leads to asymmetric, non-orthogonal interactions among velocity perturbations [22]. Also, from a mathematical perspective, our findings are extremely generic. To observe pseudo-bifurcation, all one needs to assume is that the Jacobian of the linearized system, around its equilibrium point, forms a non-normal matrix. This is a highly general condition, as non-normal matrices form a measure-one set in the space of all matrices. Our findings thus suggest that many systems previously thought to be approaching criticality may need to be reconsidered, as they could instead be stable yet non-normal systems. More generally, our results offer a different perspective on how to interpret precursory signals in dynamical systems and highlight the role of non-normal dynamics in improving our understanding and management of a wide range of physical, natural, and social systems.
@@
corresponding to the linearization of a general dynamics in the neighborhood of a stable attractor. The vector $\mathbf{x}$ quantifies the distance from the stable attractor. The matrix $\mathbf{A}$, which governs the system’s interactions, is an $n\times n$ matrix with eigenvalues $\lambda_{1},\ldots,\lambda_{n}$, all of which are assumed to have negative real parts, ensuring that the equilibrium point $\mathbf{x}=0$ remains stable. The $n$-dimensional noise term $\boldsymbol{\eta}$ is assumed to be normally distributed and $B$ quantifies the amplitude of the noise term. For simplicity, we take $B$ as scalar but reintroducing its tensorial nature does not change the main results presented below.
@@
Denote by $\mathbf{P}$ the basis transformation matrix that diagonalizes $\mathbf{A}$, that is $\mathbf{A}=\mathbf{P}\Lambda\mathbf{P}^{-1}$ with $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{n})$. Because $\mathbf{A}$ is non-normal, $\mathbf{P}$ is non-unitary, that is transforming $\mathbf{A}$ into diagonal form requires not just rotating but also scaling. Denote by $\mathbf{P}=\mathbf{U}\Sigma\mathbf{V}^{\dagger}$ the matrix $\mathbf{P}$’s singular value decomposition (SVD) and by $\mathbf{u}_{i}$ the $i$-th column vector of unitary matrix $\mathbf{U}$. Matrix $\mathbf{A}$ can be decomposed into pairwise projection operations $\mathbf{A}\sim\sum_{i,j=1}^{n}\frac{\sigma_{i}}{\sigma_{j}}\mathbf{u}_{i}\mathbf{u}_{j}^{\dagger}$ where $\sigma_{i}$ denotes the $i$-th singular value in $\Sigma=\text{diag}(\sigma_{1},\ldots,\sigma_{n})$. It is useful to consider the condition number $\kappa$ of matrix $\mathbf{A}$, defined as the ratio of the largest to the smallest singular values, $\kappa=\sigma_{1}/\sigma_{n}$, where we assume that the values are sorted without loss of generality. For a normal system, $\kappa=1$, which means that $\Sigma$ reduces to the identity matrix and the SVD of $\mathbf{P}$ reduces to $\mathbf{P}=\mathbf{U}\mathbf{V}^{\dagger}$, ensuring its unitarity and thus $\mathbf{A}^{\dagger}\mathbf{A}=\mathbf{A}\mathbf{A}^{\dagger}$. This suggests that the system’s degree of non-normality is well captured by the matrix $\mathbf{A}$’s condition number $\kappa$, as shown in the SM. Other measures of non-normality include the Kreiss constant or Henrici’s departure from normality [23].
In the remainder of this article, we assume, for simplicity, that the first $n-1$ eigenvectors are all (nearly) orthogonal to each other, with only the last one being non-orthogonal. With this assumption, $\sigma_{1}\approx\sigma_{2}\approx\ldots\approx\sigma_{n-1}\approx 1\gg\sigma_{n}=\kappa^{-1}$. This simplifies some of the subsequent analysis, but does not qualitatively change our results. As shown in the SM Appendix, so long as $\sigma_{n-1}\gg\sigma_{n}$ an assumption that is empirically supported across numerous sociotechnical and natural systems [38] - the solution of (1) exhibits non-normal transients. The SM Appendix further details that the dynamics of (1) can effectively be decomposed into two components: one along the non-normal mode $\hat{\mathbf{n}}\sim\mathbf{u}_{n}$ and one along its orthogonal reaction mode $\hat{\mathbf{r}}\sim\sum_{i=1}^{n-1}\mathbf{u}_{i}$, such that equation (1) is reduced to an effective two-dimensional system
@@
If the system’s degree of non-normality is sufficiently large, the dynamics of (1) become punctuated by transient repulsion. This can be understood from equations (2) by ignoring the stochastic term and higher-order terms. Starting at $(n,r)=(1,0)$, the variable $n$, and thus $||\mathbf{x}||_{2}$, initially decreases as the non-normal mode relaxes to zero at a rate $\alpha_{n}$ via (2a). Simultaneously, the reaction component $r$ grows rapidly at a rate $\kappa$ via (2b). When $\kappa$ is sufficiently large, to be specified below, it dominates the relaxation dynamics, leading to the growth of $||\mathbf{x}||_{2}$. This growth saturates at its peak when $\dot{r}=0$, after which the relaxation dynamics takes over, driving the system back to its stable equilibrium at $||\mathbf{x}||_{2}=0$ at rate $\lambda_{+}$ (Figure 1, bottom). We stress that this two-dimensional decomposition applies to systems of arbitrary dimension $n$.
@@
where $\theta_{t}$ is a time-dependent mean-reversion rate with memory. The exact expression of $\theta_{t}$ is given in the SM and its qualitative behavior is effectively illustrated in Figure 2 (left), capturing the key properties derived in the appendix. For $\kappa>\kappa_{c}$, where
@@
$\theta_{t}$ becomes negative over a finite time interval, corresponding to a repulsion of $r$ from its stable equilibrium at $r=0$. For systems with $\kappa>\kappa_{c}$ the stochastic dynamics of $||\mathbf{x}||_{2}$ is characterized by intermittent shifts between two regimes. For relatively small values of $||\mathbf{x}||$, the noise is dominating and the behavior appears primarily stochastic around the equilibrium point. If, by chance, the system’s trajectory is increasingly parallel to the non-normal mode, the transient amplification takes hold and the system is dominated by the approximately deterministic transient repulsion that gives rise to stochastic loops (Figure 2, right).
@@
A bifurcation in a continuous dynamical system occurs when a small change in a parameter causes a qualitative change in the stability of an equilibrium or fixed point. Mathematically, this corresponds to the real part of an eigenvalue of the equilibrium passing through zero. Formally, for a system $\dot{\mathbf{x}}=f(\mathbf{x};\mu)$, a bifurcation occurs at $(\mathbf{x}_{0},\mu_{0})$ if the Jacobian $Df_{\mathbf{x}_{0},\mu_{0}}$ has eigenvalues with zero real part. If the eigenvalue has an imaginary component, the bifurcation is classified as a Hopf bifurcation, leading to periodic limit cycles. Otherwise, it is referred to as a steady-state bifurcation, such as a saddle-node bifurcation (where two equilibria collide) or a pitchfork bifurcation (where one fixed point branches into three). These bifurcations are not just mathematical constructs but have significant real-world implications, leading to qualitatively altered system behavior. For instance, desertification in semi-arid regions, such as the Sahel, occurs when vegetation loss passes a critical threshold, leading to a rapid and often irreversible shift to desert landscapes [10]. Similarly, lake eutrophication, such as in Lake Erie, involves a bifurcation where nutrient overload causes a sudden transition from a clear-water state to a turbid, algae-dominated system [11].
@@
where $\phi\equiv\exp(\lambda_{\text{eff}})-1<0$ captures the system’s mean-reversal component with effective negative eigenvalue $\lambda_{\text{eff}}$. The system in equation (5) exhibits fluctuations with a variance equal to $B^{2}/(2|\phi|)$, and the correlation between states separated by a time interval $\Delta t$ is given by $\exp(-|\phi|\Delta t)$. Therefore, near a critical point, autocorrelation decays slowly with a characteristic time scale diverging as $\sim|\phi|^{-1}$, causing the system to relax slowly to equilibrium, a phenomenon known as critical slowing down. At the same time, the variance tends to infinity. These characteristic statistical signatures are collectively known as early-warning signals for bifurcation transitions [5, 6, 4, 7].
@@
However, while bifurcations can generate such patterns, similar dynamics can also emerge when a non-normal system experiences transient repulsion, even in the absence of an actual bifurcation. Figure 3 shows the temporal evolution of $\phi$ and $\lambda_{\text{eff}}$ obtained by fitting an Ornstein-Uhlenbeck process to the mean-field dynamics (1). During large transients, these coefficients approach zero, while the autocorrelation and variance increase in a way similar to the onset of a critical bifurcation. These observations can be explained by the fact that the transient dynamics temporarily induce a regime of repulsive, quasi-deterministic growth via (2b). At a qualitative level, the stochastic loops (Figure 2, right and Figure 3, (d)) could easily be mistaken for Hopf limit cycles. This underscores that non-normality, much like bifurcations, can generate the same early-warning signals, complicating the interpretation of these indicators. Given the similarity between these statistical patterns and those observed near critical transitions, even when the underlying non-normal system is far from or entirely devoid of true critical behavior, we refer to these transient repulsive dynamics as pseudo-bifurcations. This terminology is further justified by the fact that the natural tool to analyse the transient dynamics of non-normal operators is the pseudospectrum [23], which extends the concept of eigenvalues by replacing the condition of a vanishing characteristic determinant with the requirement that it be small, thereby capturing transient amplification effects that standard spectral analysis misses.
Distinguishing true bifurcations (if any) from pseudo-bifurcations becomes even more challenging as a system nears an actual critical bifurcation point, because, as we now show, the likelihood of observing non-normal transients increases simultaneously.
@@
Equation (4) shows that, as the system’s effective eigenvalue $\lambda_{\text{eff}}$, and thus $\lambda_{+}\lambda_{-}$, approaches zero, the critical threshold $\kappa_{c}$ at some point crosses the system’s non-normality level $\kappa$, and hence repulsive transients emerge stochastically. In other words, before entering a critical regime, the system first passes through a pseudo-critical regime (Figure 2, middle). Early warning signals that are usually used to gauge a system’s distance to criticality indicate wrongly that the system is much closer to a critical point than it actually is.
@@
In addition, complex systems are frequently governed by mechanisms such as proportional growth that drive $\mathbf{A}$ towards states of increased hierarchical organizations and hence non-normality [39, 21, 37]. The SM derives that the variance of system (5) scales like $B_{\text{eff}}\approx B\kappa^{2}/\left(\lambda_{+}\lambda_{-}(\lambda_{+}+\lambda_{-})\right)$. Similarly, the auto-correlation at lag $\Delta t$ is given by $\exp(-\Gamma(\tau)~{}\Delta t)$ with $\Gamma(\tau)\approx 1/B_{\text{eff}}$. As the system’s non-normality, and thus $\kappa$ increases, so does the system’s effective variance and auto-correlation time. This suggests an increased likelihood of false-positive markers of bifurcation in systems that dynamically self-organize into hierarchical structures.
@@
Brain connectomes - comprehensive maps of structural and functional neural connections - are essential tools for studying the brain’s complex dynamics. In the context of epilepsy, brain connectomes have been used to explore the neural underpinnings of seizures, which involve sudden, excessive synchronization of neural activity [40]. It has been hypothesized that seizures may be linked to the brain’s connectome operating near a critical state, a threshold where the brain’s dynamics are poised between order and chaos [41]. However, whether or not the brain operates at criticality during epileptic seizures remains debated [18, 17]. A mean-field model of the excitation and inhibition dynamics has been proposed to explain these phenomena [42]. This model demonstrates how non-normal transients, driven by an interplay between excitation and inhibition, can lead to large bursts of neural activity, reminiscent of seizure dynamics. Building on these observations, and because it has been shown that the connectome is strongly non-normal [21], we hypothesize that a non-normal transient burst in the brain’s dynamics may be the underlying source of the apparent increase in excitability during seizures, shifting the system into a more excitable, uncontrolled regime and revealing the inherent non-normal characteristics of the connectome.
To test this claim, we utilize an EEG recording of a patient during an epileptic seizure recorded at the Children’s Hospital Boston [43]. The measurement includes data across 23 EEG channels sampled at 256 Hz, before, during and after a seizure event that has been classified as lasting for 40 seconds (see top-left in Figure 4 for the mean-field dynamics). We first estimate the matrix $\mathbf{A}$ of (1) as $\hat{\mathbf{A}}$ obtained by fitting a VAR(1) process to the data over a rolling window of 40 seconds. In principle, we could now proceed by first determining the transition matrix $\mathbf{P}$ that diagonalizes $\hat{\mathbf{A}}$, and subsequently calculating the SVD $\mathbf{P}=\mathbf{U}\Sigma\mathbf{V}^{\dagger}$ to identify the non-normal mode $\hat{\mathbf{n}}\sim\mathbf{u}_{n}$ and reaction mode $\hat{\mathbf{r}}\sim\sum_{i=1}^{n-1}\mathbf{u}_{i}$. However, as we detail in the SM, this approach is highly sensitive to noise, especially when $\mathbf{A}$ is non-normal. Therefore, we propose an alternative approach that consists in estimating the non-normal and reaction component directly from $\hat{\mathbf{A}}$ as
where $\mathbb{S}^{n-1}$ is the unit sphere in $n$ dimensions. $\langle\mathbf{v},\hat{\mathbf{A}}\mathbf{w}\rangle$ represents the inner product (or dot product) between the vector $\mathbf{v}$ and the result of applying the matrix $\hat{\mathbf{A}}$ to the vector $\mathbf{w}$. It thus measures the degree of alignment or projection of the transformed vector $\hat{\mathbf{A}}\mathbf{w}$ onto $\mathbf{v}$. Equation (6) directly captures the intuition behind the reduced equation (2) whereby the non-normal component maximally projects onto the reaction component. From (2), we further see that $\kappa$ can be estimated directly from $\hat{\mathbf{r}}$ and $\hat{\mathbf{n}}$ via $\kappa=\sqrt{\left|\langle\hat{\mathbf{r}},\hat{\mathbf{A}}\hat{\mathbf{n}}\rangle\right/\left|\langle\hat{\mathbf{n}},\hat{\mathbf{A}}\hat{\mathbf{r}}\rangle\right|}$ thus allowing us to completely bypass the need for estimating the singular values which are numerically unstable. In the SM, we test our empirical procedure on synthetic data, finding that our procedure faithfully recovers the non-normal and reaction mode, and estimates $\kappa$ more precisely than when calculating it directly via singular value decomposition.
@@
The lower-left plot of Figure 4 reveals a marked increase in the connectome’s non-normality that occurs precisely at the onset of the epileptic seizure, persisting well beyond its clinical definition. This quantitatively demonstrates that the dynamics associated with the seizure extend far beyond the observable clinical symptoms. This is further supported by the prolonged high amplitude and variance in EEG recordings, which remain elevated long after the seizure has clinically ended.
In line with the observation that $\mathbf{A}$ becomes highly non-normal after the onset of the epileptic seizure, we observe in the top-right plots of Figure 4 that there are stochastic loops during, but not before the seizure.

### v3

@@ Introduction
Many of the dynamics driving complex systems in nature and society are marked by abrupt changes that result in transitions to fundamentally different behaviors [1, 2, 3, 4]. In recent decades, the dominant framework for understanding these transitions has centered around the concept of bifurcations [5, 6, 4, 7]. A bifurcation is a point at which a small change in a system’s control parameter causes a qualitative shift in its equilibrium state, resulting in a new and distinct configuration of system behavior. There are numerous examples of systems that display dynamics resembling those of bifurcations. For instance in ecology, predator-prey dynamics can lead to the sudden extinctions of a species [8], small shifts in climate patterns may induce inter-glacial periods [9], vegetation losses above a critical threshold results in desertification [10], or nutrient overload can cause a lake to suddenly transition from a clear-water to a algae dominated state [11]. In social systems, notable examples are financial markets that experience long periods of low volatility followed by abrupt crashes [12] and growing dissatisfaction with political conditions, which can trigger widespread unrest or revolution [13]. Health-related examples involve the sudden onset of conditions such as microbiome dysregulation [14], transitions between different states that represent distinct illnesses based on varying internal and external influences [15], depressive episodes [16], or epileptic seizures [17].
@@
As a system approaches a critical bifurcation threshold, it is often marked by increased variance and autocorrelation in observable metrics. These statistical signatures reveal that the system’s recovery from perturbations is slowing as it approaches bifurcation, alongside an increasing sensitivity to fluctuations in the system’s state–two complementary indicators that signal the proximity to a critical transition. Thus, it has been suggested that these statistical signatures serve as important early-warning signals for the onset of a looming bifurcation point [5, 6, 4, 7].
@@
Despite the insights offered by bifurcation theory, the framework itself encounters at least two major challenges. First, it relies on the existence of a well-defined control parameter and its proximity to a critical threshold. Second, it assumes the presence of a driving force that pushes the system toward or beyond this threshold. These assumptions are not only challenging to verify empirically, but also conceptually restrictive and lack general applicability. It is therefore not surprising that the question of whether certain systems are undergoing critical bifurcations continues to be a subject of debate. For example, evidence is mixed on whether the onset of epilepsy is truly driven by critical transitions [18, 17]. Similarly, it remains uncertain whether components of the Earth’s system, such as the Amazon rainforest or the Atlantic Meridional Overturning Circulation (AMOC), are approaching bifurcation points [19].
@@
In this work, we present an alternative, more general mechanism that relies solely on a stable equilibrium attractor with linear non-normal stochastic dynamics around it. We demonstrate that small perturbations around the equilibrium produce transient repulsions - we dub them pseudo-bifurcations in this context - with similar statistical signatures to those observed near genuine bifurcations. Much like the effective dimension reduction seen during the onset of a bifurcation, we find that non-normal transients are confined to movements within a two-dimensional subspace. We further find that these transients are accompanied by statistical markers such as critical slowing down and increased autocorrelation. We empirically illustrate our findings by analyzing EEG signals that capture brain activity during epileptic seizures, revealing clear evidence of non-normal dynamics. We further demonstrate that the approach of actual bifurcations is accompanied by non-normal transients, making it difficult to distinguish between the two phenomena and potentially creating a bias that suggests the system is much closer to a critical point than it actually is. Finally, we show that the onset of non-normal transients can be anticipated by measuring the system’s alignment along its non-normal mode, and provide a numerical algorithm for stable identification of these non-normal modes. This opens a new path towards the anticipation and control of instabilities in complex systems.
@@
The only requirement to observe such transients is that the system exhibits sufficient non-normality - meaning its components interact within asymmetric hierarchical structures, a feature common in many social and natural systems [20, 21]. This characteristic is also found in non-network systems, such as turbulent flows, where non-normal dynamics arise from the interplay of viscous and advective terms, which leads to asymmetric, non-orthogonal interactions among velocity perturbations [22]. From a mathematical perspective, our findings are generic. To observe pseudo-bifurcation, all one needs to assume is that the Jacobian of the linearized system, around its equilibrium point, forms a non-normal matrix. This is a general condition, as non-normal matrices form a measure-one set in the space of all matrices. Our findings thus suggest that many systems previously thought to be approaching criticality may need to be reconsidered, as they could instead be stable yet non-normal systems. More generally, our results offer a different perspective on how to interpret precursory signals in dynamical systems and highlight the role of non-normal dynamics in improving our understanding and management of a wide range of physical, natural, and social systems.
Results
@@
corresponding to the linearization of a general dynamics in the neighborhood of a stable attractor. The vector $\mathbf{x}$ quantifies the distance from the stable attractor. The matrix $\mathbf{A}$, which governs interactions of the system, is an $n\times n$ matrix with eigenvalues $\lambda_{1},\ldots,\lambda_{n}$, all of which are assumed to have negative real parts, ensuring that the equilibrium point $\mathbf{x}=0$ remains stable. The $n$-dimensional noise term $\boldsymbol{\eta}$ is assumed to be normally distributed and $B$ quantifies the amplitude of the noise term. For simplicity, we take $B$ as scalar but reintroducing its tensorial nature does not change the main results presented below.
@@
Denote by $\mathbf{P}$ the basis transformation matrix that diagonalizes $\mathbf{A}$, that is, $\mathbf{A}=\mathbf{P}\Lambda\mathbf{P}^{-1}$ with $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{n})$. Because $\mathbf{A}$ is non-normal, $\mathbf{P}$ is non-unitary, that is transforming $\mathbf{A}$ into diagonal form requires not just rotating but also scaling. Denote by $\mathbf{P}=\mathbf{U}\Sigma\mathbf{V}^{\dagger}$ the matrix $\mathbf{P}$’s singular value decomposition (SVD) and by $\mathbf{u}_{i}$ the $i$-th column vector of unitary matrix $\mathbf{U}$. The matrix $\mathbf{A}$ can be decomposed into pairwise projection operations $\mathbf{A}\sim\sum_{i,j=1}^{n}\frac{\sigma_{i}}{\sigma_{j}}\mathbf{u}_{i}\mathbf{u}_{j}^{\dagger}$ where $\sigma_{i}$ denotes the $i$-th singular value in $\Sigma=\text{diag}(\sigma_{1},\ldots,\sigma_{n})$. It is useful to consider the condition number $\kappa$ of matrix $\mathbf{A}$, defined as the ratio of the largest to the smallest singular values, $\kappa=\sigma_{1}/\sigma_{n}$, where we assume that the values are sorted without loss of generality. For a normal system, $\kappa=1$, which means that $\Sigma$ reduces to the identity matrix and the SVD of $\mathbf{P}$ reduces to $\mathbf{P}=\mathbf{U}\mathbf{V}^{\dagger}$, ensuring its unitarity and thus $\mathbf{A}^{\dagger}\mathbf{A}=\mathbf{A}\mathbf{A}^{\dagger}$. This suggests that the system’s degree of non-normality is well captured by the matrix $\mathbf{A}$’s condition number $\kappa$ (see Methods). Other measures of non-normality include the Kreiss constant or Henrici’s departure from normality [23].
In the remainder of this article, we assume, for simplicity, that the first $n-1$ eigenvectors are all (nearly) orthogonal to each other, with only the last one being non-orthogonal. With this assumption, $\sigma_{1}\approx\sigma_{2}\approx\ldots\approx\sigma_{n-1}\approx 1\gg\sigma_{n}=\kappa^{-1}$. This simplifies some of the subsequent analysis, but does not qualitatively change our results. As shown in the Method section below, so long as $\sigma_{n-1}\gg\sigma_{n}$ an assumption that is empirically supported across numerous sociotechnical and natural systems [38] - the solution of (1) exhibits non-normal transients. In that case, the dynamics of (1) can effectively be decomposed into two components: one along the non-normal mode $\hat{\mathbf{n}}\sim\mathbf{u}_{n}$ and one along its orthogonal reaction mode $\hat{\mathbf{r}}\sim\sum_{i=1}^{n-1}\mathbf{u}_{i}$, such that equation (1) is reduced to an effective two-dimensional system
@@
If the system’s degree of non-normality is sufficiently large, the dynamics of (1) becomes punctuated by transient repulsion. This can be understood from equations (2) by ignoring the stochastic term and higher-order terms. Starting at $(n,r)=(1,0)$, the variable $n$, and thus $||\mathbf{x}||_{2}$, initially decreases as the non-normal mode relaxes to zero at a rate $\alpha_{n}$ via (2a). Simultaneously, the reaction component $r$ grows rapidly at a rate $\kappa$ via (2b). When $\kappa$ is sufficiently large, as specified below, it dominates the relaxation dynamics, leading to the growth of $||\mathbf{x}||_{2}$. This growth saturates at its peak when $\dot{r}=0$, after which the relaxation dynamics takes over, restoring the system to its stable equilibrium at $||\mathbf{x}||_{2}=0$ at rate $\lambda_{+}$ (Figure 1, bottom). We stress that this two-dimensional decomposition applies to systems of arbitrary dimension $n$.
@@
where $\theta_{t}$ is a time-dependent mean-reversion rate with memory (see Methods). The qualitative behavior of $\theta_{t}$ is illustrated in Figure 2 (left). For $\kappa>\kappa_{c}$, where
@@
$\theta_{t}$ becomes negative over a finite time interval, corresponding to a repulsion of $r$ from its stable equilibrium at $r=0$. For systems with $\kappa>\kappa_{c}$ the stochastic dynamics of $||\mathbf{x}||_{2}$ is characterized by intermittent shifts between two regimes. For relatively small values of $||\mathbf{x}||$, the noise is dominating and the behavior appears primarily stochastic around the equilibrium point. If, by chance, the system’s trajectory is increasingly parallel to the non-normal mode, the transient amplification takes hold, and the system is dominated by the approximately deterministic transient repulsion that gives rise to stochastic loops (Figure 2, right).
@@
A bifurcation in a continuous dynamical system occurs when a small change in a parameter causes a qualitative change in the stability of an equilibrium or fixed point. Mathematically, this corresponds to the real part of an eigenvalue of the equilibrium passing through zero. Formally, for a system $\dot{\mathbf{x}}=f(\mathbf{x};\mu)$, a bifurcation occurs at $(\mathbf{x}_{0},\mu_{0})$ if the Jacobian $Df_{\mathbf{x}_{0},\mu_{0}}$ has eigenvalues with zero real part. If the eigenvalue has an imaginary component, the bifurcation is classified as a Hopf bifurcation, leading to periodic limit cycles. Otherwise, it is referred to as a steady-state bifurcation, such as a saddle-node bifurcation (where two equilibria collide) or a pitchfork bifurcation (where one fixed point branches into three). These bifurcations are not just mathematical constructs, but have significant real-world implications, leading to qualitatively altered system behavior. For instance, desertification in semi-arid regions, such as the Sahel, occurs when vegetation loss passes a critical threshold, leading to a rapid and often irreversible shift to desert landscapes [10]. Similarly, lake eutrophication, such as in Lake Erie, involves a bifurcation where nutrient overload causes a sudden transition from a clear-water state to a turbid, algae-dominated system [11].
@@
where $\phi\equiv\exp(\lambda_{\text{eff}})-1<0$ captures the system’s mean-reversal component with effective negative eigenvalue $\lambda_{\text{eff}}$. The system in equation (5) exhibits fluctuations with a variance equal to $B^{2}/(2|\phi|)$, and the correlation between states separated by a time interval $\Delta t$ is given by $\exp(-|\phi|\Delta t)$. Therefore, near a critical point, autocorrelation slowly decays with a characteristic time scale diverging as $\sim|\phi|^{-1}$, causing the system to relax slowly to equilibrium, a phenomenon known as critical slowing down. At the same time, the variance tends to infinity. These characteristic statistical signatures are collectively known as early-warning signals for bifurcation transitions [5, 6, 4, 7].
@@
However, while bifurcations can generate such patterns, similar dynamics can also emerge when a non-normal system experiences transient repulsion, even in the absence of an actual bifurcation. Figure 3 shows the temporal evolution of $\phi$ and $\lambda_{\text{eff}}$ obtained by fitting an Ornstein-Uhlenbeck process to the mean-field dynamics (1). During large transients, these coefficients approach zero, while the autocorrelation and variance increase in a way similar to the onset of a critical bifurcation. These observations can be explained by the fact that the transient dynamics temporarily induce a regime of repulsive, quasi-deterministic growth via (2b). At a qualitative level, the stochastic loops (Figure 2, right, and Figure 3, (d)) could easily be mistaken for Hopf limit cycles. This underscores that non-normality, much like bifurcations, can generate the same early-warning signals, complicating the interpretation of these indicators. Given the similarity between these statistical patterns and those observed near critical transitions, even when the underlying non-normal system is far from or entirely devoid of true critical behavior, we refer to these transient repulsive dynamics as pseudo-bifurcations. This terminology is further justified by the fact that the natural tool to analyze the transient dynamics of non-normal operators is the pseudospectrum [23], which extends the concept of eigenvalues by replacing the condition of a vanishing characteristic determinant with the requirement that it be small, thereby capturing transient amplification effects that standard spectral analysis misses.
@@
Distinguishing true bifurcations (if any) from pseudo-bifurcations becomes even more challenging as a system nears an actual critical bifurcation point. From equation (4) we see that as the system’s effective eigenvalue $\lambda_{\text{eff}}$, and thus $\lambda_{+}\lambda_{-}$, approaches zero, the critical threshold $\kappa_{c}$ at some point crosses the system’s non-normality level $\kappa$, and hence repulsive transients emerge stochastically. In other words, before entering a critical regime, the system first passes through a pseudo-critical regime (Figure 2, middle). Early warning signals that are usually used to gauge a system’s distance to criticality indicate wrongly that the system is much closer to a critical point than it actually is.
@@
In addition, complex systems are frequently governed by mechanisms such as proportional growth that drive $\mathbf{A}$ towards states of increased hierarchical organizations and hence non-normality [39, 21, 37]. As shown in the Methods Section below, the variance of system (5) scales like $B_{\text{eff}}\approx B\kappa^{2}/\left(\lambda_{+}\lambda_{-}(\lambda_{+}+\lambda_{-})\right)$. Similarly, the auto-correlation at lag $\Delta t$ is given by $\exp(-\Gamma(\tau)~{}\Delta t)$ with $\Gamma(\tau)\approx 1/B_{\text{eff}}$. As the system’s non-normality, and thus $\kappa$ increases, so does the system’s effective variance and auto-correlation time. This suggests an increased likelihood of false-positive markers of bifurcation in systems that dynamically self-organize into hierarchical structures.
@@
Brain connectomes, which are comprehensive maps of structural and functional neural connections, are essential tools for studying the complex dynamics of the brain. In the context of epilepsy, brain connectomes have been used to explore the neural underpinnings of seizures, which involve sudden and excessive synchronization of neural activity [40]. It has been hypothesized that seizures may be related to the brain connectome operating near a critical state, a threshold where brain dynamics is poised between order and chaos [41]. However, whether or not the brain operates at criticality during epileptic seizures remains debated [18, 17]. A mean-field model of the excitation and inhibition dynamics has been proposed to explain these phenomena [42]. This model demonstrates how non-normal transients, driven by an interplay between excitation and inhibition, can lead to large bursts of neural activity, reminiscent of seizure dynamics. Building on these observations, and because it has been shown that the connectome is strongly non-normal [21], we hypothesize that a non-normal transient burst in the brain’s dynamics may be the underlying source of the apparent increase in excitability during seizures, shifting the system into a more excitable, uncontrolled regime and revealing the inherent non-normal characteristics of the connectome.
@@
To test this claim, we utilize an EEG recording of a patient during an epileptic seizure recorded at the Children’s Hospital Boston [43]. The measurement includes data across 23 EEG channels sampled at 256 Hz, before, during and after a seizure event that has been classified as lasting for 40 seconds (see top-left in Figure 4 for the mean-field dynamics). We empirically estimate the system’s non-normality $\kappa$, as well as non-normal and reaction mode (see Methods). The lower left plot of Figure 4 reveals a marked increase in the connectome’s non-normality that occurs precisely at the onset of the epileptic seizure, persisting well beyond its clinical definition. This quantitatively demonstrates that the dynamics associated with the seizure extend far beyond the observable clinical symptoms. This is further supported by the prolonged high amplitude and variance in EEG recordings, which remain elevated long after the seizure has clinically ended. In line with the observation that $\mathbf{A}$ becomes highly non-normal after the onset of the epileptic seizure, we observe in the top right plots of Figure 4 that there are stochastic loops during, but not before the seizure.